OpenStudy (anonymous):

what is the derivative given y=sin^3(3x)+(cos^2)2x

OpenStudy (anonymous):

Well, I'll write all the basic formulas you need to do this... Derivatives of the following are: x^n=n*x^(n-1) sinx=cosx cosx=-sinx Use the chain rule..

OpenStudy (anonymous):

First, assign sinx as p,q or any variable you want..

OpenStudy (anonymous):

And cosx as some particular variable..

OpenStudy (anonymous):

Then, use the first formula..i.e, x^n=n*x^(n-1)..

OpenStudy (anonymous):

Are you there??

OpenStudy (anonymous):

Well, anyway..multiply whatever you get with the derivative of sinx and cosx.. In this case...there's another number multiplied to the x... Suppose it's sin5x..then the derivative would be... cos5x*5 or 5cos5x... I hope you understand...

OpenStudy (anonymous):

I think you can work it out on your own now..